OpenStudy (anonymous):

order from least to greatest: -0.7, 0.4, 1/4, 1/7

OpenStudy (paki):

-0.7, 1/7, 1/4, 0.4

OpenStudy (anonymous):

Thank you! That's what I thought, but I wasn't sure.

OpenStudy (paki):

welcome :-)

OpenStudy (anonymous):

Can you help me with one more? Just check my answer? Order 2/7, 1/5, and 3/4 from least to greatest.

OpenStudy (anonymous):

My answer: 2/7, 1/5, 2/4

OpenStudy (paki):

ok, let me check

OpenStudy (paki):

1/5, 2/7, 2/4... 1/5= 0.2 2/7= 0.28 2/4= 0.5

OpenStudy (anonymous):

It iS 3/4

OpenStudy (paki):

are you saying about the last one

OpenStudy (paki):

if it is the last one... then the order will be like this... 1/5. 2/7, 3/4 :-)

OpenStudy (anonymous):

Okay, thanks!!

OpenStudy (paki):

pleasure